OpenStudy (anonymous):

HELP! MEDALS WILL BE GIVEN I PROMISE! What is the standard form of an ellipse?

OpenStudy (anonymous):

i promise i will give medals.

OpenStudy (anonymous):

You need to complete the squares on the LHS: 32x^2 - 256x + 9y^2 + 72y = -368 32(x^2 - 8x + _) + 9(y^2 + 8y + _) = -368 +32( _) + 9(_) Note the factors on the RHS to keep the equation balanced 32(x^2 - 8x + 16) + 9(y^2 + 8y + 16) = -368 + 32(16) + 9(16) (16 = (8/4)^2, the technique for completing the square) 32(x - 4)^2 + 9(y + 4)^2 = 288 now divide by 288 to put in standard form for ellipses: (x - 4)^2 / 9 + (y + 4)^2 / 32 = 1 ellipse with center at (4,-4), vertical major axis

OpenStudy (anonymous):

yaaaay thanls so much medal coming